# Gradient Flow Finite Element Discretisations with Energy-Based $hp$–Adaptivity for the Gross-Pitaevskii Equation with Angular Momentum Rotation

### v1

@@ 2.3.2. Energy-based orthogonal projection
$\nabla_{a_{u}}\mathsf{E}(u)=u\qquad\text{on }\mathbb{H}.$ (23)
We note that we could omit taking the real-part on the left-hand side of (21), if instead of employing the real Gâteaux derivative $\mathsf{E}^{\prime}$ of $\mathsf{E}$ at $u$, the Wirtinger derivative were considered. This, however, would render the ensuing analysis unnecessarily more technical.
@@
Remark 2.5.
For any $u\in\mathbb{H}$, we note that
$u(\bm{\mathsf{x}})\overline{\mathsf{G}(u(\bm{\mathsf{x}}))}=|u(\bm{\mathsf{x}})||\mathsf{G}(u(\bm{\mathsf{x}}))|\qquad\text{for a.e. }\bm{\mathsf{x}}\in\Omega.$ (28)
This follows from an argument analogous to [28, Lemma 5.5]. Specifically, from (27), we observe that $\mathsf{G}(u)$ is the unique minimiser of the real-valued functional
$\mathsf{H}_{u}(w):=\frac{1}{2}\Re\left(a_{u}(w,w)\right)-\Re\left((u,w)_{\mathrm{L}^{2}(\Omega)}\right),\qquad w\in\mathbb{H}.$
Hence, if we let $\widetilde{w}:=|\mathsf{G}(u)|\mathrm{e}^{i\varphi_{u}}$, where $\varphi_{u}\in[0,2\pi)$ is the phase of $u$ (i.e., $u=|u|\mathrm{e}^{i\varphi_{u}}$), then a straightforward calculation reveals that $\mathsf{H}_{u}(\widetilde{w})\leq\mathsf{H}_{u}(\mathsf{G}(u))$. This implies that $\widetilde{w}=\mathsf{G}(u)$ holds almost everywhere in $\Omega$, which directly yields (28).
@@ 2.3.3. Projected gradient flow
for some initial guess $u^{0}\in\mathcal{S}_{\mathbb{H}}$, where we have employed the identity (23) in the second equality. Using the representation (25), the dynamical system (29) is given by
@@
where we define
$\gamma_{u(t)}$ $=\frac{\|u(t)\|^{2}_{\mathrm{L}^{2}(\Omega)}}{(\mathsf{G}(u(t)),u(t))_{\mathrm{L}^{2}(\Omega)}}.$ (30b)
for some initial $u(0)=u^{0}\in\mathcal{S}_{\mathbb{H}}$. We note that both the numerator as well as the denominator in (30b) are real-valued; in addition, taking into account (28), we observe that
$\gamma_{u(t)}>0\qquad\text{if }u(t)\neq 0,$ (31)
and there is no change of phase throughout the flow.
In the following result, we discuss some properties of the flow defined by the dynamical system (30). To this end, we assume that $u(t)$ is well-defined for all $t\geq 0$.
Theorem 2.6.
Given that the flow $u(t)$ is defined for all $t\geq 0$, the following properties hold:
  1. (a)
    If $u(0)=u_{0}\in\mathcal{S}_{\mathbb{H}}$, then $u(t)\in\mathcal{S}_{\mathbb{H}}$ for all $t\geq 0$;
  2. (b)
    The energy $\mathsf{E}(t):=\mathsf{E}(u(t))$ is monotone decreasing in $t\geq 0$;
  3. (c)
    If $u_{0}\in\mathcal{S}_{\mathbb{H}}$, then the flow $u(t)$ converges to an eigenfunction $u^{\star}\in\mathcal{S}_{\mathbb{H}}$ of the Gross–Pitaevskii equation (13) associated with the eigenvalue given by
    $\lambda^{\star}=\frac{1}{(\mathsf{G}_{u^{\star}}(u^{\star}),u^{\star})_{\mathrm{L}^{2}(\Omega)}}.$ (32)
Proof of Theorem 2.6 (a).
Due to the orthogonal projection onto the tangent space $\mathbb{T}_{u}$, the flow from (30) is $\mathrm{L}^{2}(\Omega)$-norm preserving. Indeed, since
$\dot{u}(t)=-\mathsf{P}_{u(t)}(u(t))\in\mathbb{T}_{u(t)},$ (33)
we find by a straightforward calculation that
@@
which yields the first claim. ∎
Proof of Theorem 2.6 (b).
Taking the (real) derivative of the function
$g(t):=\mathsf{E}(u(t)),\qquad t\geq 0,$ (34)
and invoking (22) implies that $g^{\prime}(t)=\langle\mathsf{E}^{\prime}(u),\dot{u}\rangle=\Re\left(a_{u}(u,\dot{u})\right)$, where we have suppressed the dependence of $u=u(t)$ on $t$ for better readability. Combining (24) and (30), we deduce that
$g^{\prime}(t)=\Re(a_{u}(\mathsf{P}_{u}(u),\dot{u}))=-\Re(a_{u}(\dot{u},\dot{u}))=-a_{u}(\dot{u},\dot{u})\leq 0,$ (35)
cf. Proposition 2.2; i.e., the energy $\mathsf{E}(u(t))$ is monotonically decreasing. This establishes (b). ∎
In the remaining part (c) of the proof, we may assume without loss of generality that the function $g$ from (34) satisfies $g^{\prime}(t)\neq 0$ for all $t\geq 0$. Otherwise, if there exists $t_{\star}\geq 0$ such that $g^{\prime}(t_{\star})=0$, then $u(t_{\star})$ is an eigenfunction of the GPE (12). Indeed, from (35) we deduce that $\dot{u}(t_{\star})=0$, which, upon recalling (30) and (31), leads to $u(t_{\star})=\lambda\mathsf{G}(u(t_{\star}))$ for $\lambda:=\Gamma(u(t_{\star}))>0$. In view of (26) we arrive at
$a_{u(t_{\star})}(u(t_{\star}),v)=\lambda a_{u(t_{\star})}(\mathsf{G}(u(t_{\star})),v)=\lambda(u(t_{\star}),v)_{\mathrm{L}^{2}(\Omega)},$
for any $v\in\mathbb{H}$. The proof of Theorem 2.6 (c) requires the following auxiliary result.
Lemma 2.7.
Suppose that the solution $u$ of (33) is $C^{2}$ in time. Then, the function $g$ from (34) satisfies the bound $g^{\prime\prime}(t)\geq-2g^{\prime}(t)$ for all $t>0$.
Proof.
For time-dependent $\mathbb{H}$-valued functions $u,v,w$, a straightforward calculation shows that
$\frac{{\mathsf{d}}}{\,\mathsf{d}t}a_{u}(v,w)=a_{u}(\dot{v},w)+a_{u}(v,\dot{w})+4\beta\int_{\Omega}\Re\left(u\overline{\dot{u}}\right)v\overline{w}\,\mathsf{d}\bm{\mathsf{x}},$ (36)
and therefore
$\frac{\mathsf{d}}{\,\mathsf{d}t}\Re(a_{u}(v,w))=\Re(a_{u}(\dot{v},w))+\Re(a_{u}(v,\dot{w}))+4\beta\int_{\Omega}\Re\left(u\overline{\dot{u}}\right)\Re\left(v\overline{w}\right)\,\mathsf{d}\bm{\mathsf{x}}.$
Moreover, recalling (24) and (30a), for any $v\in\mathbb{T}_{u}$, where $u$ is the solution of (33), the following holds
$\Re(a_{u}(u,v))=\Re(a_{u}(\mathsf{P}_{u}(u),v))=-\Re(a_{u}(\dot{u},v)).$
Taking the temporal derivative yields
$\frac{\mathsf{d}}{\,\mathsf{d}t}\Re(a_{u}(u,v))=-\frac{\mathsf{d}}{\,\mathsf{d}t}\Re(a_{u}(\dot{u},v)),$
which, upon using (36), results in the identity
$\Re(a_{u}(\dot{u},v))+4\beta\int_{\Omega}\Re\left(u\overline{\dot{u}}\right)\Re\left(u\overline{v}\right)\,\mathsf{d}\bm{\mathsf{x}}=-\Re(a_{u}(\ddot{u},v))-4\beta\int_{\Omega}\Re\left(u\overline{\dot{u}}\right)\Re\left(\dot{u}\overline{v}\right)\,\mathsf{d}\bm{\mathsf{x}},$
for any $v\in\mathbb{T}_{u}$ (independent) of time. Then, for any given $t>0$, inserting $v=\dot{u}(t)\in\mathbb{T}_{u(t)}$, cf. (33), we deduce the equality
$\Re(a_{u}(\dot{u},\dot{u}))+4\beta\left\|\Re\left(u\overline{\dot{u}}\right)\right\|^{2}_{\mathrm{L}^{2}(\Omega)}=-\Re(a_{u}(\ddot{u},\dot{u}))-4\beta\int_{\Omega}\Re\left(u\overline{\dot{u}}\right)\left|\dot{u}\right|^{2}\,\mathsf{d}\bm{\mathsf{x}}.$
Applying (35), and rearranging terms, we deduce that
$\Re(a_{u}(\ddot{u},\dot{u}))=g^{\prime}(t)-4\beta\left\|\Re\left(u\overline{\dot{u}}\right)\right\|^{2}_{\mathrm{L}^{2}(\Omega)}-4\beta\int_{\Omega}\Re\left(u\overline{\dot{u}}\right)\left|\dot{u}\right|^{2}\,\mathsf{d}\bm{\mathsf{x}}.$
Furthermore, again with the aid of (35) and (36), we have
$g^{\prime\prime}(t)=-\frac{\mathsf{d}}{\,\mathsf{d}t}\Re(a_{u}(\dot{u},\dot{u}))$ $=-2\Re(a_{u}(\ddot{u},\dot{u}))-4\beta\int_{\Omega}\Re\left(u\overline{\dot{u}}\right)\left|\dot{u}\right|^{2}\,\mathsf{d}\bm{\mathsf{x}}.$
By combining the above two equations it follows that
$g^{\prime\prime}(t)=-2g^{\prime}(t)+8\beta\left\|\Re\left(u\overline{\dot{u}}\right)\right\|^{2}_{\mathrm{L}^{2}(\Omega)}+4\beta\int_{\Omega}\Re\left(u\overline{\dot{u}}\right)\left|\dot{u}\right|^{2}\,\mathsf{d}\bm{\mathsf{x}}.$ (37)
In light of (30a), and invoking (31) and (28), we observe that
$u\overline{\dot{u}}=u\left(-\overline{u}+\gamma_{u}\overline{\mathsf{G}(u)}\right)=-|u|^{2}+\gamma_{u}|u||\mathsf{G}(u)|$
is a real-valued function. Hence, we conclude that
$2\Re\left(u\overline{\dot{u}}\right)^{2}+\Re\left(u\overline{\dot{u}}\right)\left|\dot{u}\right|^{2}=|\dot{u}|^{2}\left(2|u|^{2}+u\overline{\dot{u}}\right)=|\dot{u}|^{2}|u|\left(|u|+\gamma_{u}|\mathsf{G}(u)|\right)\geq 0$
Exploiting this bound in (37) completes the proof. ∎
We are now prepared to complete the proof of Theorem 2.6.
Proof of Theorem 2.6 (c).
Based on Lemma 2.7, we can apply the argument pursued in the proof of [34, Thm. 7.1] to conclude that the limit $u(t)\to u^{\star}$ as $t\to\infty$ exists. Moreover, upon taking the limit in (30a), we find that $u^{\star}=\lambda^{\star}\mathsf{G}(u^{\star})$, where
$\lambda^{\star}=\frac{\|u^{\star}\|^{2}_{\mathrm{L}^{2}(\Omega)}}{(\mathsf{G}(u^{\star}),u^{\star})_{\mathrm{L}^{2}(\Omega)}}=\frac{1}{(\mathsf{G}(u^{\star}),u^{\star})_{\mathrm{L}^{2}(\Omega)}}>0,$
cf. (31). Therefore, by exploiting (26), we infer that
$\Re\left(a_{u^{\star}}(u^{\star},v)\right)=\lambda^{\star}\Re\left(a_{u^{\star}}(\mathsf{G}(u^{\star}),v)\right)=\lambda^{\star}\Re\left((u^{\star},v)_{\mathrm{L}^{2}(\Omega)}\right)\qquad\forall v\in\mathbb{H}.$
It now follows from Remark 2.4 that $u^{\star}\not\equiv 0$ is an eigenfunction of the GPE (12) with corresponding eigenvalue $\lambda^{\star}$. ∎
@@ 2.4. Discrete gradient flow
For the purpose of computing an approximation of the continuous gradient flow trajectory from (30), we employ the forward Euler time stepping method. Specifically, for a given initial value $u^{0}\in\mathcal{S}_{\mathbb{H}}$ this yields a sequence $\{u^{n}\}_{n\geq 0}\subset\mathcal{S}_{\mathbb{H}}$ which, for $n\geq 0$, is defined by
@@
$0<\tau_{\min}\leq\tau_{n}\leq\tau_{\max}<2\qquad\forall n\geq 0$ (39)
The scheme (38) is called discrete gradient flow iteration (GFI). We note that this iteration scheme coincides with the one from [29], and, up to an additional term in the weighted inner-product that incorporates the angular momentum, resembles the GFI proposed in [28, §4]. In particular, we may borrow the following result from [29, Thm. 3.3] in a modified form, which can also (largely) be obtained by following along the lines of [28, §4] (with some minor adaptations); we provide a rough sketch for the latter approach in Appendix B.
@@
For a sufficiently small lower bound $\tau_{\min}$, there exists a constant $\tau_{\min}<\tau_{\max}<2$ such that, for any sequence of time-steps $\{\tau_{n}\}_{n}\subset[\tau_{\min},\tau_{\max}]$, the sequence $\{u^{n}\}_{n}$ generated by the GFI (38a) satisfies the following properties:
@@
    We have that $\mathsf{E}(u^{n+1})\leq\mathsf{E}(u^{n})$ for all $n\geq 0$. Indeed, there exists a constant $C_{\tau}>0$ (depending on $\tau_{\min}$ and $\tau_{\max}$) such that
    $\mathsf{E}(u^{n+1})-\mathsf{E}(u^{n})\geq C_{\tau}\left\|u^{n+1}-u^{n}\right\|_{\mathrm{L}^{2}(\Omega)}^{2}$ (40)
@@
    The limit $\mathsf{E}^{\star}_{\infty}:=\lim_{n\to\infty}\mathsf{E}(u^{n})$ is well-defined. Moreover, there exists a strongly convergent subsequence $\{u^{n_{k}}\}_{k}$ in $\mathbb{H}$ with a limit $u^{\star}_{\infty}\in\mathcal{S}_{\mathbb{H}}$, and $\mathsf{E}(u^{\star}_{\infty})=\mathsf{E}^{\star}_{\infty}$.
@@
    Furthermore, for the limit $u^{\star}_{\infty}$ from (b), upon defining $\lambda^{\star}_{\infty}$ as in (32), we have that
@@
    $a_{u^{\star}_{\infty}}(u^{\star}_{\infty},v)=\lambda^{\star}_{\infty}(u^{\star}_{\infty},v)_{\mathrm{L}^{2}(\Omega)}\qquad\forall v\in\mathbb{H},$
    i.e. $(\lambda^{\star}_{\infty},u^{\star}_{\infty})$ is a solution to the Gross–Pitaevskii EVP (13).
3. $hp$–Adaptive Gradient Flow Finite Element Discretisation

### v2

@@ 2.3.2. Energy-based orthogonal projection
[…]
We note that we could omit taking the real-part on the left-hand side of (23), if, instead of employing the real Gâteaux derivative $\mathsf{E}^{\prime}$ of $\mathsf{E}$ at $u$, the Wirtinger derivative were considered. This, however, would render the ensuing analysis unnecessarily more technical.
@@ 2.3.3. Projected gradient flow
for some initial guess $u^{0}\in\mathcal{S}_{\mathbb{H}}$, where we have employed the identity (25) in the second equality. Using the representation (27), the above dynamical system (30) can also be expressed by
@@
with
$\gamma_{u(t)}=\frac{\|u(t)\|^{2}_{\mathrm{L}^{2}(\Omega)}}{(\mathsf{G}(u(t)),u(t))_{\mathrm{L}^{2}(\Omega)}},$
for some initial $u(0)=u^{0}\in\mathcal{S}_{\mathbb{H}}$. In light of (28) and the coercivity of $a_{u(t)}$, cf. Proposition 2.3, we observe that
$(\mathsf{G}(u(t)),u(t))_{\mathrm{L}^{2}(\Omega)}=\overline{(u(t),\mathsf{G}(u(t)))}_{\mathrm{L}^{2}(\Omega)}=a_{u(t)}(\mathsf{G}(u(t)),\mathsf{G}(u(t)))>0;$
i.e., $\gamma_{u(t)}>0$ for any $t>0$. Furthermore, since $\dot{u}(t)=-\mathsf{P}_{u(t)}(u(t))\in\mathbb{T}_{u(t)}$, a straightforward calculation yields that
@@
i.e. the flow preserves the $\mathrm{L}^{2}(\Omega)$-norm. Consequently, for $u_{0}\in\mathcal{S}_{\mathbb{H}}$, we conclude that
$\gamma_{u(t)}=\frac{1}{(\mathsf{G}(u(t)),u(t))_{\mathrm{L}^{2}(\Omega)}}>0.$ (31b)
Finally, we remark that the energy $\mathsf{E}$ is monotone decreasing along the trajectory $u(t)$ as $t$ progresses. Indeed, letting $g(t):=\mathsf{E}(u(t))$, $t\geq 0$, we invoke (24) to obtain
$g^{\prime}(t)=\langle\mathsf{E}^{\prime}(u(t)),\dot{u}(t)\rangle=\Re\left(a_{u(t)}(u(t),\dot{u}(t))\right).$
Then, combining (26) and (31), we deduce that
$g^{\prime}=\Re(a_{u}(\mathsf{P}_{u}(u),\dot{u}))=-\Re(a_{u}(\dot{u},\dot{u}))=-a_{u}(\dot{u},\dot{u})\leq 0,$
cf. Proposition 2.3; i.e., the energy $\mathsf{E}(u(t))$ is monotonically decreasing in $t$.
@@ 2.4. Discrete gradient flow
For the purpose of computing an approximation of the continuous projected Sobolev gradient flow trajectory from (31), we employ the forward Euler time stepping method. Specifically, for a given initial value $u^{0}\in\mathcal{S}_{\mathbb{H}}$, this yields a sequence $\{u^{n}\}_{n\geq 0}\subset\mathcal{S}_{\mathbb{H}}$, which, for $n\geq 0$, is defined by
@@
$0<\tau_{\min}\leq\tau_{n}\leq\tau_{\max}<2\qquad\forall n\geq 0.$ (33)
The scheme (32) is called discrete gradient flow iteration (GFI). We note that this iteration coincides with the one from [30], and, up to an additional term in the weighted inner-product that incorporates the angular momentum, resembles the GFI proposed in [29, §4]. In particular, we may borrow the following result from [30, Thm. 3.4] in a modified form, which can also (largely) be obtained by following along the lines of [29, §4] (with some minor adaptations).
@@
For a sufficiently small lower bound $\tau_{\min}$, there exists a constant $\tau_{\max}\in(\tau_{\min},2)$ such that, for any sequence of time-steps $\{\tau_{n}\}_{n}\subset[\tau_{\min},\tau_{\max}]$, the resulting sequence $\{u^{n}\}_{n}$ generated by the GFI (32) satisfies the following properties:
@@
    We have that $\mathsf{E}(u^{n+1})\leq\mathsf{E}(u^{n})$ for all $n\geq 0$. More precisely, there exists a constant $C_{\tau}>0$ (only depending on $\tau_{\min}$ and $\tau_{\max}$) such that
    $\mathsf{E}(u^{n})-\mathsf{E}(u^{n+1})\geq C_{\tau}\left\|u^{n+1}-u^{n}\right\|_{\mathrm{L}^{2}(\Omega)}^{2}$ (34)
@@
    The limit $\mathsf{E}^{\star}:=\lim_{n\to\infty}\mathsf{E}(u^{n})$ is well-defined. Moreover, there exists a strongly convergent subsequence $\{u^{n_{k}}\}_{k}$ in $\mathbb{H}$ with a limit $u^{\star}\in\mathcal{S}_{\mathbb{H}}$, and $\mathsf{E}(u^{\star})=\mathsf{E}^{\star}$.
@@
    Furthermore, for the limit $u^{\star}$ from (b), and upon letting
    $\lambda^{\star}:=\frac{1}{(\mathsf{G}_{u^{\star}}(u^{\star}),u^{\star})_{\mathrm{L}^{2}(\Omega)}},$
    we have that
    $a_{u^{\star}}(u^{\star},v)=\lambda^{\star}(u^{\star},v)_{\mathrm{L}^{2}(\Omega)}\qquad\forall v\in\mathbb{H},$ (35)
    i.e. $(\lambda^{\star},u^{\star})$ is a (weak) solution to the Gross–Pitaevskii EVP (13).
Proof.
We will sketch the proof by focussing on the required modifications of the techniques presented in [29, §4] for the more general case including angular momentum. We proceed in several steps.
  1. (i)
    A straightforward calculation reveals that the mass, without the scaling (32a), grows in the iteration step (32b). To this end, we can follow (exactly) along the lines of[29, §4.1]. In particular, upon exploiting (32b), i.e.,
    $\widehat{u}^{n+1}-u^{n}=-\tau_{n}\mathsf{P}_{u^{n}}(u^{n})=-\tau_{n}u^{n}+\frac{\tau_{n}}{(\mathsf{G}(u^{n}),u^{n})_{\mathrm{L}^{2}(\Omega)}}\mathsf{G}(u^{n}),$
    we obtain that $(u^{n},\widehat{u}^{n+1}-u^{n})_{\mathrm{L}^{2}(\Omega)}=0$, and thus
    $\left\|\widehat{u}^{n+1}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}=\left\|u^{n}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}+\left\|\widehat{u}^{n+1}-u^{n}\right\|_{\mathrm{L}^{2}(\Omega)}^{2}\geq\left\|u^{n}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}.$ (36)
  2. (ii)
    Next, we will verify the (discrete) energy dissipation. This is done in several steps and will be summarised in the following. Since the energy, in addition to the angular momentum term, is slightly different from [29], we will provide some more details. Firstly, using (28) and applying a similar argument as in (i), we observe that $a_{u^{n}}(\mathsf{G}(u^{n}),\widehat{u}^{n+1}-u^{n})=0,$ and
    $a_{u^{n}}(\widehat{u}^{n+1}-u^{n},\widehat{u}^{n+1}-u^{n})=-\tau_{n}a_{u^{n}}(u^{n},\widehat{u}^{n+1}-u^{n}).$ (37)
    Equivalently,
    $a_{u^{n}}(u^{n},u^{n})=a_{u^{n}}(\widehat{u}^{n+1},\widehat{u}^{n+1})+\left(\nicefrac{{2}}{{\tau_{n}}}-1\right)a_{u^{n}}(\widehat{u}^{n+1}-u^{n},\widehat{u}^{n+1}-u^{n}),$ (38)
    cf. the proof of [29, Lem. 4.5]. Moreover, invoking (1) and (14), we find that
    $a_{u^{n}}(u^{n},u^{n})=2\mathsf{E}(u^{n})+\beta\int_{\Omega}|u^{n}|^{4}\,\mathsf{d}\bm{\mathsf{x}},$
    and similarly
    $a_{u^{n}}(\widehat{u}^{n+1},\widehat{u}^{n+1})=2\mathsf{E}(\widehat{u}^{n+1})+2\beta\int_{\Omega}|u^{n}|^{2}|\widehat{u}^{n+1}|^{2}\,\mathsf{d}\bm{\mathsf{x}}-\beta\int_{\Omega}|\widehat{u}^{n+1}|^{4}\,\mathsf{d}\bm{\mathsf{x}}.$ (39)
    Combining the identities (38) and (39), we deduce that
    $\mathsf{E}(u^{n})$ $-\mathsf{E}(\widehat{u}^{n+1})$
    $=-\frac{\beta}{2}\int_{\Omega}\left(|u^{n}|^{2}-|\widehat{u}^{n+1}|^{2}\right)^{2}\,\mathsf{d}\bm{\mathsf{x}}+\left(\frac{1}{\tau_{n}}-\frac{1}{2}\right)a_{u^{n}}(\widehat{u}^{n+1}-u^{n},\widehat{u}^{n+1}-u^{n}).$
    Furthermore, repeating the calculations in the proof of Theorem [29, Lem. 4.5] (with $\beta$ replaced by $2\beta$), we arrive at
    $\mathsf{E}(u^{n})-\mathsf{E}(\widehat{u}^{n+1})$ $\geq\left(\frac{1}{\tau_{n}}-\frac{1}{2}\right)a_{0}(\widehat{u}^{n+1}-u^{n},\widehat{u}^{n+1}-u^{n})-\beta\int_{\Omega}|\widehat{u}^{n+1}-u^{n}|^{4}\,\mathsf{d}\bm{\mathsf{x}}$
    $\quad+2\beta\left(\frac{1}{\tau_{n}}-\frac{5}{2}\right)\int_{\Omega}|\widehat{u}^{n+1}-u^{n}|^{2}|u^{n}|^{2}\,\mathsf{d}\bm{\mathsf{x}}.$
    In turn, for $\tau_{n}\leq\nicefrac{{2}}{{5}}$, we have that
    $\mathsf{E}(u^{n})-\mathsf{E}(\widehat{u}^{n+1})\geq\left(\frac{1}{\tau_{n}}-\frac{1}{2}\right)a_{0}(\widehat{u}^{n+1}-u^{n},\widehat{u}^{n+1}-u^{n})-\beta\int_{\Omega}|\widehat{u}^{n+1}-u^{n}|^{4}\,\mathsf{d}\bm{\mathsf{x}}.$ (40)
  3. (iii)
    Based on (36), (37) and (40), it can be shown by an inductive argument, cf. the proof of [29, Lem. 4.7], that there exists $0<\tau_{\max}\leq\nicefrac{{2}}{{5}}$ (depending on $\mathsf{E}(u^{0})$ and $\beta$) such that the energy decays. Specifically, similarly to [29, Eq. (26)], we derive the bound
    $\mathsf{E}(u^{n})-\mathsf{E}(\widehat{u}^{n+1})\geq c\,a_{0}(\widehat{u}^{n+1}-u^{n},\widehat{u}^{n+1}-u^{n})\geq 0,$ (41)
    for some $c>0$ independent of $n$ (but depending on $\tau_{\max}$ and $\beta$). Then,
    $\mathsf{E}(u^{n+1})\leq\mathsf{E}(\widehat{u}^{n+1})\leq\mathsf{E}(u^{n}),$ (42)
    where the first inequality follows from the mass growth obtained in (i). This proves the first part of (a). Concerning the bound (34), we simply refer to the proof in [30, Thm 3.4].
  4. (iv)
    By the energy decay (42) and the non-negativity of $\mathsf{E}$, cf. (6), it immediately follows that the sequence $\{\mathsf{E}(u^{n})\}_{n}$ has a limit $\mathsf{E}^{\star}\geq 0$. In turn, by virtue of the coercivity bound (22), the sequence $\{u^{n}\}_{n}$ is uniformly bounded in $\mathbb{H}$, and thus has a weakly convergent subsequence $\{u^{n_{j}}\}_{j}$ in $\mathbb{H}$ with a limit $u^{\star}\in\mathbb{H}$; thanks to the Rellich-Kondrachov compactness theorem we may assume that the convergence is strong in $\mathrm{L}^{p}(\Omega;\mathbb{C})$, for $1\leq p<6$ in dimensions $d\in\{2,3\}$; in particular, we have $u^{\star}\in\mathcal{S}_{\mathbb{H}}$. Moreover, using (41) and (42), and employing Proposition 2.3, we obtain the bounds
    $\mathsf{E}(u^{n})-\mathsf{E}(u^{n+1})\geq\mathsf{E}(u^{n})-\mathsf{E}(\widehat{u}^{n+1})\geq\widetilde{c}\left\|\widehat{u}^{n+1}-u^{n}\right\|^{2}_{\mathbb{H}},$
    for a constant $\widetilde{c}>0$. Hence, from the convergence of the energy, we further deduce that $\left\|\widehat{u}^{n+1}-u^{n}\right\|_{\mathbb{H}}\to 0$. Then, by repeating the calculations in [29, Thm. 4.9], we find that
@@
    $\nicefrac{{1}}{{\lambda^{n_{j}}}}:=(\mathsf{G}(u^{n_{j}}),u^{n_{j}})_{\mathrm{L}^{2}(\Omega)}\to(\mathsf{G}(u^{\star}),u^{\star})_{\mathrm{L}^{2}(\Omega)}=:\nicefrac{{1}}{{\lambda^{\star}}}\quad\text{as}\ j\to\infty,$ (43)
    from which (35) can be obtained. In particular, the claim in (c) is verified.
  5. (v)
    Next, we will show that $\mathsf{E}(u^{\star})=\mathsf{E}^{\star}$. Our proof is slightly different from the one in [29]. First of all, using the definitions of the energy functional and of the weighted inner-product from (1) and (14), respectively, we find that
    $2\left|\mathsf{E}(u^{\star})-\mathsf{E}(u^{n_{j}})\right|$ $=\left|a_{u^{\star}}(u^{\star},u^{\star})-\beta\int_{\Omega}|u^{\star}|^{4}\,\mathsf{d}\bm{\mathsf{x}}-a_{u^{n_{j}}}(u^{n_{j}},u^{n_{j}})+\beta\int_{\Omega}|u^{n_{j}}|^{4}\,\mathsf{d}\bm{\mathsf{x}}\right|$
    $\leq|\lambda^{\star}-\lambda^{n_{j}}|+|\lambda^{n_{j}}-a_{u^{n_{j}}}(u^{n_{j}},u^{n_{j}})|+\beta\left|\left\|u^{\star}\right\|_{\mathrm{L}^{4}(\Omega)}^{4}-\left\|u^{n_{j}}\right\|_{\mathrm{L}^{4}(\Omega)}^{4}\right|;$
    the first term vanishes in the limit $j\to\infty$ thanks to (43), while the last term vanishes as $j\to\infty$ due to the strong convergence in $\mathrm{L}^{4}(\Omega)$. Concerning the second summand, in light of (32b), we first observe that
    $\frac{1}{\tau_{n_{j}}}a_{u^{n_{j}}}(\widehat{u}^{n_{j}+1},u^{n_{j}})-\frac{1-\tau_{n_{j}}}{\tau_{n_{j}}}a_{u^{n_{j}}}(u^{n_{j}},u^{n_{j}})=\lambda^{n_{j}},$
    and in turn
    $|\lambda^{n_{j}}-a_{u^{n_{j}}}(u^{n_{j}},u^{n_{j}})|=\frac{1}{\tau_{n_{j}}}\left|a_{u^{n_{j}}}(\widehat{u}^{n_{j}+1}-u^{n_{j}},u^{n_{j}})\right|.$
    Making use of (37) yields
    $|\lambda^{n_{j}}-a_{u^{n_{j}}}(u^{n_{j}},u^{n_{j}})|$ $=\tau^{-2}_{n_{j}}a_{u^{n_{j}}}(\widehat{u}^{n_{j}+1}-u^{n_{j}},\widehat{u}^{n_{j}+1}-u^{n_{j}})\lesssim\tau_{\min}^{-2}\left\|\widehat{u}^{n_{j}+1}-u^{n_{j}}\right\|^{2}_{\mathbb{H}};$
    here, we employed Proposition 2.3 and the uniform boundedness of the sequence $\{u^{n_{j}}\}_{j}$. Furthermore, recalling the strong convergence $\left\|\widehat{u}^{n+1}-u^{n}\right\|_{\mathbb{H}}\to 0$ as $n\to\infty$ from (iv), we conclude that the second summand vanishes as well. Hence, $\lim_{j\to\infty}\left|\mathsf{E}(u^{\star})-\mathsf{E}(u^{n_{j}})\right|=0$.
  6. (vi)
    It remains to verify that the subsequence $\{u^{n_{j}}\}_{j}$ converges to $u^{\star}$ strongly in $\mathbb{H}$. Since the proof of this step in the present work is more involved than in the case without angular momentum, we shall provide full details. From (iv) and (v), we recall that $\mathsf{E}(u^{n_{j}})\to\mathsf{E}(u^{\star})$ and $u^{n_{j}}\to u^{\star}$ strongly in $\mathrm{L}^{p}(\Omega)$, for $1\leq p<6$, and $u^{n_{j}}\rightharpoonup u^{\star}$ weakly in $\mathbb{H}$. Thereby, this implies that
    $\lim_{j\to\infty}\left(\frac{1}{2}\left\|u^{n_{j}}\right\|^{2}_{\mathbb{H}}-\int_{\Omega}i\omega\overline{u^{n_{j}}}(\bm{\mathsf{A}}\cdot\nabla u^{n_{j}})\,\mathsf{d}\bm{\mathsf{x}}\right)=\frac{1}{2}\left\|u^{\star}\right\|^{2}_{\mathbb{H}}-\int_{\Omega}i\omega\overline{u^{\star}}(\bm{\mathsf{A}}\cdot\nabla u^{\star})\,\mathsf{d}\bm{\mathsf{x}}.$
    In particular, if we can show that
    $\left|\int_{\Omega}\overline{u^{n_{j}}}(\bm{\mathsf{A}}\cdot\nabla u^{n_{j}})\,\mathsf{d}\bm{\mathsf{x}}-\int_{\Omega}\overline{u^{\star}}(\bm{\mathsf{A}}\cdot\nabla u^{\star})\,\mathsf{d}\bm{\mathsf{x}}\right|$
    $\leq\int_{\Omega}\left|u^{n_{j}}-u^{\star}\right|\left|\bm{\mathsf{A}}\cdot\nabla u^{n_{j}}\right|\,\mathsf{d}\bm{\mathsf{x}}+\left|\int_{\Omega}\left(\overline{u^{\star}}(\bm{\mathsf{A}}\cdot\nabla u^{n_{j}})-\overline{u^{\star}}(\bm{\mathsf{A}}\cdot\nabla u^{\star})\right)\,\mathsf{d}\bm{\mathsf{x}}\right|$
    $=:I_{1}+I_{2}.$
    For the first term, we recall that $\{u^{n_{j}}\}$ is a weakly convergent sequence in $\mathbb{H}$, and thus is uniformly bounded. Therefore, the Cauchy–Schwarz inequality yields that
    for some constant $C>0$ independent of $j$. To deal with the second term, we note that $v\mapsto\int_{\Omega}\overline{u^{\star}}(\bm{\mathsf{A}}\cdot\nabla v)\,\mathsf{d}\bm{\mathsf{x}}$ is a bounded linear operator on $\mathbb{H}$, and thus $I_{2}\to 0$ as $j\to\infty$ thanks to the weak convergence of the sequence.
This completes the proof. ∎
3. hphp–Adaptive Gradient Flow Finite Element Discretisation